OpenStudy (kkrpbra):

A red lionfish is about 18 inches long. Some islands in the Bahamas have reported a tenfold increase in these fish in the last year. Scientists saw a lionfish eat 20 small fish in just 30 minutes. Study the numbers in this paragraph. The square root of one of the numbers is 3√2. Which number is it?

OpenStudy (anonymous):

what the heck kind of question is this? study the numbers in the paragraph and find the square root of them? yikes

OpenStudy (anonymous):

answer is 18, because \[\sqrt{18}=\sqrt{9\times 2}=\sqrt{9}\times \sqrt{2}=3\sqrt{2}\]

OpenStudy (anonymous):

i have to copy this down to show how silly math questions can get . is this from a book?

OpenStudy (kkrpbra):

kidbiz3000.com